OpenStudy (anonymous):

2(x + y) = 2x + 2y is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

jimthompson5910 (jim_thompson5910):

have a look at this page http://www.quia.com/jg/332021list.html if you're completely stuck

OpenStudy (anonymous):

still suck lol

jimthompson5910 (jim_thompson5910):

go back to that page I gave you and try replacing numbers with variables

jimthompson5910 (jim_thompson5910):

for example, one line gives you Commutative Property of Addition (Numbers) 3 + 7 = 7 + 3 replace 7 with x and replace 3 with y to get x+y = y+x

jimthompson5910 (jim_thompson5910):

So in general, the Commutative Property of Addition says that x+y = y+x is true for any real numbers x and y